OpenStudy (anonymous):

How do I solve with a calculator: Sin(theta)=0.25

OpenStudy (anonymous):

arcsin(0.25)

OpenStudy (anonymous):

the same as sin^(-1)

OpenStudy (anonymous):

the answer I get is .25 the answer is says is correct is 14.5 and 165.5???Im confused

OpenStudy (unklerhaukus):

\[\sin(\theta)=0.25\] \[\theta=\arcsin\left(0.25\right)\] [asin] [ ( ] [0] [.] [2] [5] [ ) ] [ = ]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=sin^-1%280.25%29 The answer in degrees is 14.5

OpenStudy (anonymous):

The answer you get is in radians, which is also correct

OpenStudy (anonymous):

I get the 14.5 answer but what about the 165.5?

OpenStudy (anonymous):

Think of it as a measurement of a triangle, you can go both the minus way and the positive way it doesn't matter. So if you get the angle 14.5 degrees there also the angle -14.5 which can be expressed as 180-14.5 = 165.5

OpenStudy (anonymous):

ahh, thats awesome, all I needed to know thank you

OpenStudy (anonymous):

No problem! :)